### v1

# Learning Task-adaptive Quasi-stiffness Control for a Powered Transfemoral Prosthesis

Abstract
While significant advancements have been made in the mechanical and task-specific controller designs of powered transfemoral prostheses, developing a task-adaptive control framework that generalizes across various locomotion modes and terrain conditions remains an open problem. This study proposes a task-adaptive learning quasi-stiffness control framework for powered prostheses that generalizes across tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part. Quasi-stiffness is defined as the slope of the human joint’s torque-angle relationship. To accurately obtain the torque-angle relationship in a new task, a Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in the task. Then a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of a new task from multiple human demonstrations and estimated target features. Based on the torque-angle relationship of the new task, a quasi-stiffness control approach is designed for a powered prosthesis. Finally, the proposed framework is validated through practical examples, including varying speed and incline walking tasks. The proposed framework has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
I Introduction
Transfemoral amputees suffer ambulation difficulties when performing demanding locomotion modes, such as climbing ramps and stairs. Powered transfemoral prostheses have the potential to restore the lost function of the joints and enhance the mobility of the amputees [1, 2]. While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8, 9, 10] and controller designs [11, 12, 13, 14, 15, 16, 17] of powered prostheses over the past decades, developing a prosthesis control framework that generalizes across a variety of locomotion modes and terrain conditions remains a challenge [2].
Impedance-based control is the most commonly used approach in powered prostheses because of its ability to reproduce natural walking. These methods segment the gait cycle into several sub-phases, each performing an impedance control law. Previous researchers manually tuned the impedance parameters for each task, such as flat walking [3], ramp climbing [18], and stair climbing [19]. However, the heuristic parameter tuning procedure for a specific task takes at least several hours [20]. As shown in Fig. 1, transfemoral amputees may perform different tasks (locomotion modes and terrain conditions) in daily walking. The time required for tuning a powered prosthesis for each task and each patient presents a major barrier to its clinical feasibility [21]. To reduce the tuning time, reinforcement learning (RL) is implemented in [14, 22] to automatically tune the impedance parameters. Nevertheless, these approaches are limited to flat walking task and knee joint only at present. Recent studies have achieved prostheses handling variable-task walking. Data-driven, impedance-based controllers are respectively designed for varying speeds and inclines walking [20], stair climbing [23], and sit/stand [24]. An adaptive stair controller enables transfemoral amputee patients to climb stairs of varying stair heights [25]. These controllers are still required to extend to a unified framework for variable-task walking. To date, developing powered transfemoral prostheses controllers that generalize across tasks remains an open problem.
Quasi-stiffness control is a subset of the impedance-based control strategies[2]. This strategy determines the impedance parameters of each sub-phase from the torque-angle relationship of human’s lower limb joints. The slope of the torque-angle relationship is called ”quasi-stiffness” [2, 26]. The desired prosthesis torque is then determined by the quasi-stiffness and equilibrium angle as . The torque–angle relationships can be modified for quasi-stiffness control of powered prostheses in different tasks. For example, the stair controller modulates the torque-angle relationship based on the stair height for stair climbing tasks. In addition, it is reported that quasi-stiffness control strategies have the potential to improve gait symmetry and achieve speed-adaption without subject-specific or speed-specific tuning [21]. Inspired by the quasi-stiffness control strategy, this study proposes a learning task-adaptive quasi-stiffness control framework that generalizes across a variety of tasks. Two issues are presented and need to be resolved. The first one is to automatically construct joints’ torque-angle relationship in a new task. The second one is to design quasi-stiffness controllers based on the generated torque-angle relationship.
The torque-angle relationship is determined by the joint angle and torque and is widely used in the design and control of powered exoskeleton [27, 28] and prosthesis [21, 25]. The peak values (named as target features in this study) in the profiles of joint kinematics and kinetics vary with tasks and are often used to represent the joint kinematics and kinetics during walking tasks [14, 22]. However, most torque-angle relationships and target features are obtained in the laboratory by a motion capture system. It is impossible to collect joint kinematics and kinetics for all tasks. This study proposed an imitation learning-based method that can reconstruct the torque-angle relationships of new tasks by learning from multiple demonstrated torque-angle relationships. For example, reconstructing a relationship at a new walking speed from learning the existing demonstrated relationship at different speeds (or a new relationship at a new ramp inclination, etc.). To reconstruct the torque-angle relationship of a new task, the Gaussian Process Regression (GPR) [29, 30] is introduced to predict the target features of the joint angles and torques based on the demonstrated features and task variables (such as speed and terrain parameters). Furthermore, the new task torque-angle relationship can be reconstructed by the target features and human-demonstrated torque-angle relationships. While many fitting methods can use these target features for torque-angle relationship reconstruction, such as polynomial interpolation [31] and Fourier Series fitting [32], these methods consider only trajectory fitting instead of the multiple demonstrated relationships. Instead, learning from multiple demonstrations methods such as probabilistic trajectory methods can be introduced. Gaussian Mixture Model/Regression (GMM-GMR) [33, 34] was applied to reproduce joint angles and learning trajectory distribution through learning from the demonstrated database. However, this method makes it difficult to modify the learned trajectories to fulfill additional new constraints, such as passing through specific target features. A Kernelized Movement Primitives (KMP) [35] method was presented that allows the reproduced learning from demonstration trajectories passing through the specific via-points. Inspired by this, KMP is employed to construct the new torque-angle relationship from multiple demonstrated relationships and target features. After obtaining the torque-angle relationship, the quasi-stuffiness control strategy is introduced to design a task-adaptive controller that generalizes across tasks.
The main contributions are summarized as follows:
1.
A learning-based torque-angle relationship construction method is proposed for various tasks.
2.
A quasi-stiffness control strategy based on the learned torque-angle relationship is presented for powered transfemoral prostheses.
3.
Examples of varying speed and varying incline walking using the proposed quasi-stiffness control framework are conducted and validated through walking experiments with a transfemoral amputee.
II Method
As shown in Fig. 2, the proposed task-adaptive learning quasi-stiffness control framework consists of a learning from human demonstrations part and a quasi-stiffness control part. The learning-based part is composed of two models: a GPR-based target features prediction model and a KMP-based task-specific torque-angle relationship reconstruction model.
II-A Task Estimation
Our previous studies and some other studies presented various task estimation methods, such as vision-based methods [36, 37] for terrain condition estimation and IMU (inertial measurement unit)-based methods [20, 38, 39] for speed estimation. These methods can be applied to estimate the task parameter (where denotes the walking speed, represents the terrain parameters) for the proposed framework.
II-B GPR-based Target Features Prediction
An open-source dataset [40] is implemented as the human demonstrations, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights). To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the demonstration method for target features prediction, as shown in Fig. 2(b). Considering the limited human demonstration size, GPR, a nonparametric kernel-based probabilistic model, adeptly balances fitting accuracy with model complexity, mitigating the risk of overfitting in small-size training samples. Assume there are demonstrations and define the train data as
where denotes the number of the features; is the task parameter of the -th demonstration (-th task). is the output features (joint angles/torques) of the -th demonstration. For the task , the GPR [29, 30] is intended to transform the input task vector into a output value by , where represents the Gaussian noise satisfying . The training dataset can be described by the GPR model as
where represents the mean function and is set to zero [30]; denotes the covariance function (also called Gaussian kernel) and can be selected in [29]. The target features’ prior distribution in satisfies
where is the covariance matrix. To generate prediction for a new task input , the predicted function is following a joint Gaussian distribution:
The posterior distribution of the prediction at a new input is given by
The predicted mean value and corresponding variance of the conditional distribution can be given by [30]
where , denotes the covariance matrix between training and new input, and denotes the covariance between the new input vector.
Thus, the new output is predicted as the mean values. By employing the GPR method, the new task’s target features including joint angles and joint torques can be represented as where represents the number of target features.
II-C KMP-based Torque-angle relationship Reconstruction
After obtaining the kinematics and kinetics target features of the new task, the KMP is employed to reconstruct the torque-angle relationship with the target features of the new task, as shown in Fig. 2(c). Denote the set of demonstrated torque-anlge relationship by where represents the joint kinematics and is the joint kinetics. and denote the length of the torque-angle relationship and the number of the demonstrations. GMM [34] is introduced to encode the demonstrated dataset and estimate the relationship probability distribution
where , , and represent prior probability, mean, and covariance, respectively. is the number of the Gaussian component. Furthermore, a probabilistic reference relationship can be retrieved by GMR
where and represent the reference torque and angle respectively, and denote mean and variance respectively. As mentioned before, we expect that the new torque-angle relationship can pass through the target features of the new task. Therefore, KMP is employed to learn from demonstrations and fulfill the additional constraints. Considering a parametric torque-angle relationship can be formulated by KMP as
where is a -dimensional basis function and is defined in [35]. is the weight vector and can be assumed to be normally distributed following . Thus, the probability distribution of the parametric torque-angle relationship stratifies
The new torque-angle relationship is expected to pass through the target features. The target features is assumed satisfying and
Finally, KMP requires matching the reference torque-angle relationship with the parametric torque-angle relationship and the target features. Inspired by [35, 41, 42], to obtain and , the Kullback-Leibler (KL) divergence is introduced to design the cost function as
where denotes the KL divergence between the two probability distributions. A detailed solution to minimization of the cost function can be found in [35]. By employing the KMP, the torque-angle relationship of the new task can be reconstructed and applied in the quasi-stiffness control.
II-D Quasi-stiffness Estimation and Controller Design
For the powered prostheses, the desired torque determined by the ”quasi-stiffness” control law is defined as
where denotes the quasi-stiffness, represents the equilibrium angle. It is noted that the torque-angle relationship during the stance phase is approximately linear [26, 43]. Inspired by this, as shown in Fig. 2(d), this study applies a linear regression to estimate the quasi-stiffness of the torque-angle relationship, and the equilibrium angle is equal to the value of the intersection point between the regression equation and the x-axis. For the prosthesis control, we divide the gait phase into four sub-phases: stance flexion, stance extension, swing flexion, and swing extension. The sub-phase switching follows a finite-state machine (FSM), as shown in Fig. 3. For the stance phase, the torque-angle relationship is approximately linear, which can be approximated by a linear equation. While the swing phase is much more complex than the stance phase [43], we choose two linear parts that can achieve swing flexion and extension and use linear functions to approximate these two parts.
III Results
III-A Experimental Setup
The powered transfemoral prosthesis used in this study is shown in Fig. 6(a). Both knee and ankle joints are driven by frameless motors (ILM motor kit, RoboDrive, Seefeld, Germany) with a 50:1 harmonic reducer (CSD-25-50-2A-GR, Nagano Prefecture, Japan). The frameless motors are actuated by drives (G-MOLTWIR50/100SEO, Elmo Motion Control, Petah Tikva, Israel). A 6-axis force/torque sensor (Sunrise Instruments, Nanning, China) is employed to measure the (ground reaction forces) GRFs.
A transfemoral amputee patient is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework. The participant provided informed consent, and the Southern University of Science and Technology Institutional Review Board approved the experiments.
III-B Experimental Results of the Proposed Framework
Due to the space limitations, this study presents an example of varying walking speed tasks to show how the learning quasi-stiffness control framework works. The learning process of the GPR-based target feature estimation and KMP-based torque-angle relationship construction is offline on the joint kinematics and kinetics obtained from the open-source dataset [40].
1) Target Features Estimation: Fig. 2(b) shows the experimental results of target features prediction, where the blue dots denote the demonstrated target features and the purple dots represent the predicted target features of a new task (new speed). The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks.
2) Torque-angle relationship Reconstruction: Fig. 4 shows the torque-angle relationship learning and reconstruction in a new task. The distribution of the demonstrated torque-angle relationship can be learned by GMM. Based on the learned GMM, the reference torque-angle relationship can be reproduced by GMR. As shown in Fig. 4(a), the blue torque-angle curves denote the human demonstrations of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR. As shown in Fig. 4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the demonstrated torque-angle relationships (blue curves) and target features (purple dots) in a new task.
3) Quasi-stiffness Estimation: Fig. 5 shows using linear functions to approximate several parts of the knee and ankle torque-angle relationships in the varying speed walking tasks. Table I shows the estimated stiffness and equilibrium angles from the torque-angle relationships of different speed tasks and these impedance parameters are used in the powered transfemoral prosthesis experiments.
4) Performance of the Proposed Method in Experiments: Fig. 6(b) shows a series of photos of a transfemoral amputee performing varying speed and inclines walking tasks. Fig. 7 presents that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which also indicate that the quai-stiffness can improve gait symmetry. Fig. 8 shows the root mean squared error (RMSE) in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments. The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry.
| Task (m/s) | Joint | Quasi-stiffness (Nm/deg) | Equilibrium Angle (deg) | ||||||
| 0.4 | knee | 4.13 | 0.59 | 0.08 | 0.19 | 12 | 22 | 30 | 15 |
| ankle | 0.82 | 6.14 | 0.53 | 0.59 | -5 | -2 | 8 | -8 | |
| 0.6 | knee | 5.68 | 0.70 | 0.17 | 0.26 | 12 | 21 | 57 | 17 |
| ankle | 0.76 | 5.05 | 0.53 | 0.64 | -5 | 0 | 8 | -8 | |
| 0.8 | knee | 5.79 | 0.79 | 0.27 | 0.45 | 13 | 22 | 69 | 18 |
| ankle | 0.78 | 4.23 | 0.45 | 0.69 | -6 | -4 | 9 | -7 | |
| 1.0 | knee | 5.85 | 0.85 | 0.36 | 0.52 | 13 | 22 | 76 | 20 |
| ankle | 1.03 | 3.89 | 0.36 | 0.72 | -5 | -7 | 10 | -8 | |
IV Discussion
IV-A Evaluation of the Proposed Control Method
This work proposed a learning task-adaptive quasi-stiffness controller for a powered transfemoral prosthesis that generalizes across tasks. The proposed method has the potential to quickly and autonomously configure the impedance parameters for variable tasks. Experimental results and walking performance (see the accompanying video) highlight its capacity to replicate able-bodied biomechanics across variable tasks. As shown in Fig. 7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking speed and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which indicates that the quai-stiffness control has the potential to improve gait symmetry. Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
IV-B Limitations and Future Work
A limitation of this study is that the torque-angle relationship of the swing phase tends to be more nonlinear and complex as the task becomes more difficult (higher walking speed or more challenging terrain conditions), which means that using linear regression to estimate quasi-stiffness may introduce larger errors. In future work, we will investigate the use of nonlinear regression to estimate the swing phase quasi-stiffness. Nevertheless, this work focuses on designing a powered transfemoral prosthesis control framework that generalizes across various tasks. Experimental results and walking performance (see the accompanying video) preliminarily verify that the framework has the potential to achieve task-adaptive control. Future studies will invite more participants to verify the performance of the proposed framework in outdoor experiments.
V Conclusion
This study proposed a learning task-adaptive quasi-stiffness control framework for a powered transfemoral prosthesis working over variable tasks. Learning from the human demonstration method is developed for the torque-angle relationship reconstruction in a new task, which consists of a GPR-based target features prediction model and a KMP-based trajectories reconstruction model. A transfemoral amputee participant demonstrated the control framework’s ability to generalize across tasks by varying speeds and inclines walking experiments. The experimental results validate that the proposed control framework has the potential to reproduce biomimetic joint torques and angles for a powered prosthesis during variable-task walking.

### v2

@@
-Learning Task-adaptive Quasi-stiffness Control for a Powered Transfemoral Prosthesis
+Tuning-free Quasi-stiffness Control Framework of a Powered Transfemoral Prosthesis for Task-adaptive Walking
 Abstract
-While significant advancements have been made in the mechanical and task-specific controller designs of powered transfemoral prostheses, developing a task-adaptive control framework that generalizes across various locomotion modes and terrain conditions remains an open problem.
-This study proposes a task-adaptive learning quasi-stiffness control framework for powered prostheses that generalizes across tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part.
-Quasi-stiffness is defined as the slope of the human joint’s torque-angle relationship.
-To accurately obtain the torque-angle relationship in a new task, a Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in the task.
-Then a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of a new task from multiple human demonstrations and estimated target features.
+Impedance-based control represents a prevalent strategy in the development of powered transfemoral prostheses.
+However, creating a task-adaptive, tuning-free controller that effectively generalizes across diverse locomotion modes and terrain conditions continues to be a significant challenge.
+This letter proposes a tuning-free and task-adaptive quasi-stiffness control framework for powered prostheses that generalizes across various walking tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part.
+A Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in a new task.
+Subsequently, a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of the new task from multiple human reference trajectories and estimated target features.
 Based on the torque-angle relationship of the new task, a quasi-stiffness control approach is designed for a powered prosthesis.
-Finally, the proposed framework is validated through practical examples, including varying speed and incline walking tasks.
-The proposed framework has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
-I Introduction
-Transfemoral amputees suffer ambulation difficulties when performing demanding locomotion modes, such as climbing ramps and stairs.
-Powered transfemoral prostheses have the potential to restore the lost function of the joints and enhance the mobility of the amputees [1, 2].
-While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8, 9, 10] and controller designs [11, 12, 13, 14, 15, 16, 17] of powered prostheses over the past decades, developing a prosthesis control framework that generalizes across a variety of locomotion modes and terrain conditions remains a challenge [2].
+Finally, the proposed framework is validated through practical examples, including varying speeds and inclines walking tasks.
+Notably, the proposed framework not only aligns with but frequently surpasses the performance of a benchmark finite state machine impedance controller (FSMIC) without necessitating manual impedance tuning and has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
+Index Terms:
+Tuning-free, task-adaptive, prosthesisI Introduction
+Individuals with transfemoral amputations suffer difficulties in ambulation, particularly during demanding locomotion activities such as ascending ramps and stairs.
+Powered transfemoral prostheses hold promise in restoring the lost joint function and significantly enhancing the mobility of these individuals [1, 2].
+While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8] and controller designs [9, 10, 11, 12, 13, 14, 15, 16] of powered prostheses over the past y, developing a prosthesis control framework that generalizes across various locomotion modes and terrain conditions remains a challenge [2].
 Impedance-based control is the most commonly used approach in powered prostheses because of its ability to reproduce natural walking.
@@
 However, the heuristic parameter tuning procedure for a specific task takes at least several hours [19].
-As shown in Fig.
-1, transfemoral amputees may perform different tasks (locomotion modes and terrain conditions) in daily walking.
-The time required for tuning a powered prosthesis for each task and each patient presents a major barrier to its clinical feasibility [21].
-To reduce the tuning time, reinforcement learning (RL) is implemented in [14, 22] to automatically tune the impedance parameters.
+Transfemoral amputees may perform different tasks (locomotion modes and terrain conditions) in daily walking.
+The time required to tune a powered prosthesis for each task and each patient presents a significant barrier to its clinical feasibility [20].
+To reduce the tuning time, reinforcement learning (RL) is implemented in [12, 21] to tune the impedance parameters automatically.
 Nevertheless, these approaches are limited to flat walking task and knee joint only at present.
 Recent studies have achieved prostheses handling variable-task walking.
-Data-driven, impedance-based controllers are respectively designed for varying speeds and inclines walking [20], stair climbing [23], and sit/stand [24].
-An adaptive stair controller enables transfemoral amputee patients to climb stairs of varying stair heights [25].
+Data-driven, impedance-based controllers are respectively designed for varying speeds and inclines walking [19], stair climbing [22], and sitting/standing [23].
+An adaptive stair controller enables transfemoral amputee patients to climb stairs of varying heights [24].
 These controllers are still required to extend to a unified framework for variable-task walking.
@@
 Quasi-stiffness control is a subset of the impedance-based control strategies [2].
-This strategy determines the impedance parameters of each sub-phase from the torque-angle relationship of human’s lower limb joints.
+While traditional impedance control requires tuning numerous impedance parameters to generate biologically joint torques, this quasi-stiffness strategy determines the impedance parameters of each sub-phase from the torque-angle relationship of human lower limb joints.
 The slope of the torque-angle relationship is called ”quasi-stiffness” [2, 25].
@@
 The torque–angle relationships can be modified for quasi-stiffness control of powered prostheses in different tasks.
-For example, the stair controller modulates the torque-angle relationship based on the stair height for stair climbing tasks.
-In addition, it is reported that quasi-stiffness control strategies have the potential to improve gait symmetry and achieve speed-adaption without subject-specific or speed-specific tuning [21].
-Inspired by the quasi-stiffness control strategy, this study proposes a learning task-adaptive quasi-stiffness control framework that generalizes across a variety of tasks.
-Two issues are presented and need to be resolved.
-The first one is to automatically construct joints’ torque-angle relationship in a new task.
-The second one is to design quasi-stiffness controllers based on the generated torque-angle relationship.
-The torque-angle relationship is determined by the joint angle and torque and is widely used in the design and control of powered exoskeleton [27, 28] and prosthesis [21, 25].
-The peak values (named as target features in this study) in the profiles of joint kinematics and kinetics vary with tasks and are often used to represent the joint kinematics and kinetics during walking tasks [14, 22].
+For example, the stair controller modulates the torque-angle relationship based on the stair height for stair climbing tasks [24].
+While quasi-stiffness control strategies have been proven can improve gait symmetry and achieve speed-adaption without speed-specific tuning [20], how to adapt to other tasks automatically is still a challenge.
+Inspired by the previous research on quasi-stiffness control strategies, this study proposes a tuning-free quasi-stiffness control framework that generalizes across a variety of tasks.
+Firstly, a novel task-adaptive torque-angle relationship reconstruction method is proposed for autonomously generating the torque-angle relationship of a new task.
+The torque-angle relationship is constructed by the the joint angle and torque and its profile depends strongly on the peak values of the joint angle and torque.
+The peak values (named as target features in this study) in the profiles of joint angles and torques vary with tasks and are often used to represent the joint kinematics and kinetics during walking tasks [12, 21].
 However, most torque-angle relationships and target features are obtained in the laboratory by a motion capture system.
-It is impossible to collect joint kinematics and kinetics for all tasks.
-This study proposed an imitation learning-based method that can reconstruct the torque-angle relationships of new tasks by learning from multiple demonstrated torque-angle relationships.
-For example, reconstructing a relationship at a new walking speed from learning the existing demonstrated relationship at different speeds (or a new relationship at a new ramp inclination, etc.).
-To reconstruct the torque-angle relationship of a new task, the Gaussian Process Regression (GPR) [29, 30] is introduced to predict the target features of the joint angles and torques based on the demonstrated features and task variables (such as speed and terrain parameters).
-Furthermore, the new task torque-angle relationship can be reconstructed by the target features and human-demonstrated torque-angle relationships.
-While many fitting methods can use these target features for torque-angle relationship reconstruction, such as polynomial interpolation [31] and Fourier Series fitting [32], these methods consider only trajectory fitting instead of the multiple demonstrated relationships.
-Instead, learning from multiple demonstrations methods such as probabilistic trajectory methods can be introduced.
-Gaussian Mixture Model/Regression (GMM-GMR) [33, 34] was applied to reproduce joint angles and learning trajectory distribution through learning from the demonstrated database.
-However, this method makes it difficult to modify the learned trajectories to fulfill additional new constraints, such as passing through specific target features.
-A Kernelized Movement Primitives (KMP) [35] method was presented that allows the reproduced learning from demonstration trajectories passing through the specific via-points.
-Inspired by this, KMP is employed to construct the new torque-angle relationship from multiple demonstrated relationships and target features.
-After obtaining the torque-angle relationship, the quasi-stuffiness control strategy is introduced to design a task-adaptive controller that generalizes across tasks.
+It is only possible to collect joint kinematics and kinetics for some tasks.
+This study proposes an imitation learning-based method that can reconstruct the torque-angle relationships of new tasks by learning from multiple reference torque-angle relationships.
+For example, reconstructing a relationship at a new walking speed from learning the existing reference relationships at different speeds (or a new relationship at a new ramp inclination, etc.).
+To reconstruct the torque-angle relationship of a new task, the Gaussian Process Regression (GPR) [26, 27] is introduced to predict the target features of the joint angles and torques based on the reference features and task variables (such as speeds and terrain parameters).
+Considering the limited human reference size, GPR, a nonparametric kernel-based probabilistic model, adeptly balances fitting accuracy with model complexity, mitigating the risk of over-fitting in small-size training samples.
+Furthermore, the torque-angle relationship of the new task can be reconstructed by the target features and human reference torque-angle relationships.
+While many fitting methods can use these target features for torque-angle relationship reconstruction, such as polynomial interpolation [28] and Fourier Series fitting [29], these methods consider only trajectory fitting instead of the multiple reference relationships.
+Instead, learning from multiple reference trajectory methods, such as probabilistic trajectory methods can be introduced.
+Gaussian Mixture Model/Regression (GMM/GMR) [30, 31] was applied to reproduce joint angles and learning trajectory distribution through learning from the reference database.
+However, this method makes it difficult to modify the learned trajectories to fulfill additional constraints, such as passing through specific target features.
+A Kernelized Movement Primitives (KMP) [32] method was presented that allows the reproduced learning from reference trajectories passing through the specific via-points.
+Inspired by this, KMP is employed to construct the new torque-angle relationship from multiple reference relationships and target features.
+Secondly, the quasi-stuffiness control strategy is introduced after obtaining the torque-angle relationship to design a task-adaptive controller that generalizes across tasks.
 The main contributions are summarized as follows:
 1.
-A learning-based torque-angle relationship construction method is proposed for various tasks.
+A learning-based task-adaptive torque-angle relationship construction method is proposed for powered transfemoral prostheses, which can autonomously generate each task-specific torque-angle relationship during continuously varying tasks.
+A tuning-free quasi-stiffness control method based on the learned torque-angle relationship is presented for powered transfemoral prostheses.
 2.
-A quasi-stiffness control strategy based on the learned torque-angle relationship is presented for powered transfemoral prostheses.
-3.
-Examples of varying speed and varying incline walking using the proposed quasi-stiffness control framework are conducted and validated through walking experiments with a transfemoral amputee.
+Experimental results of varying speed and varying incline walking tasks show that the proposed control framework reproduces biomimetic kinematics and kinetics similar to able-bodied references, meets the performance of the most widely used (one of the state of the art) FSMIC without requiring manual impedance tuning, and has the potential to enable autonomous task adaptation during continuously varying walking tasks.
 II Method
+II-A Overview of the Framework
 As shown in Fig.
-2, the proposed task-adaptive learning quasi-stiffness control framework consists of a learning from human demonstrations part and a quasi-stiffness control part.
-The learning-based part is composed of two models: a GPR-based target features prediction model and a KMP-based task-specific torque-angle relationship reconstruction model.
-II-A Task Estimation
-Our previous studies and some other studies presented various task estimation methods, such as vision-based methods [36, 37] for terrain condition estimation and IMU (inertial measurement unit)-based methods [20, 38, 39] for speed estimation.
-These methods can be applied to estimate the task parameter (where denotes the walking speed, represents the terrain parameters) for the proposed framework.
-II-B GPR-based Target Features Prediction
-An open-source dataset [40] is implemented as the human demonstrations, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights).
-To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the demonstration method for target features prediction, as shown in Fig.
-2(b).
-Considering the limited human demonstration size, GPR, a nonparametric kernel-based probabilistic model, adeptly balances fitting accuracy with model complexity, mitigating the risk of overfitting in small-size training samples.
-Assume there are demonstrations and define the train data as
-where denotes the number of the features; is the task parameter of the -th demonstration (-th task).
-is the output features (joint angles/torques) of the -th demonstration.
+1, the proposed tuning-free, task-adaptive quasi-stiffness control framework consists of a task estimation part, a torque-angle relationship reconstruction part, and a quasi-stiffness control part.
+The task estimation part perceives amputee locomotion information and terrain conditions.
+It provides task parameters ( where denotes walking speed and represents terrain parameters, such as inclination and stair height .) for the next part.
+Previous research, including our own, has explored a variety of task estimation methodologies, encompassing vision-based approaches [33, 34] for assessing terrain conditions and IMU (Inertial Measurement Unit)-based strategies [19, 35, 36] for estimating walking speed.
+These methodologies facilitate the estimation of task parameters .
+The torque-angle relationship reconstruction part includes a GPR-based target features prediction model and a KMP-based task-specific torque-angle relationship reconstruction model.
+This part automatically constructs the joints’ torque-angle relationship in a new task.
+The quasi-stiffness control part determines the parameters and applies to a quasi-stiffness controller based on the generated torque-angle relationship.
+The details of these two parts are as follows.
+II-B Target Features Prediction
+An open-source dataset [37] is implemented as the human reference trajectories, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights).
+To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the reference trajectory method for target features prediction, as shown in Fig.
+1(b).
+Assume there are reference trajectories and define the train data as
+where denotes the number of the features; is the task parameter of the -th reference trajectory (-th task).
+is the output features (joint angles/torques) of the -th reference trajectory.
 For the task , the GPR [26, 27] is intended to transform the input task vector into a output value by , where represents the Gaussian noise satisfying .
@@
 By employing the GPR method, the new task’s target features including joint angles and joint torques can be represented as where represents the number of target features.
-II-C KMP-based Torque-angle relationship Reconstruction
+II-C Torque-angle Relationship Reconstruction
 After obtaining the kinematics and kinetics target features of the new task, the KMP is employed to reconstruct the torque-angle relationship with the target features of the new task, as shown in Fig.
-2(c).
-Denote the set of demonstrated torque-anlge relationship by where represents the joint kinematics and is the joint kinetics.
-and denote the length of the torque-angle relationship and the number of the demonstrations.
-GMM [34] is introduced to encode the demonstrated dataset and estimate the relationship probability distribution
+1(c).
+Denote the set of reference torque-angle relationship by where represents the joint kinematics and is the joint kinetics.
+and denote the length of the torque-angle relationship and the number of the reference trajectories.
+GMM [31] is introduced to encode the reference dataset and estimate the relationship probability distribution
 where , , and represent prior probability, mean, and covariance, respectively. is the number of the Gaussian component. Furthermore, a probabilistic reference relationship can be retrieved by GMR
@@
 As mentioned before, we expect that the new torque-angle relationship can pass through the target features of the new task.
-Therefore, KMP is employed to learn from demonstrations and fulfill the additional constraints.
+Therefore, KMP is employed to learn from reference trajectores and fulfill the additional constraints.
 Considering a parametric torque-angle relationship can be formulated by KMP as
@@
 where denotes the quasi-stiffness, represents the equilibrium angle.
+For the prosthesis control, following the quasi-stiffness trends observed in able-bodied subjects [20], we divide the gait phase into four sub-phases: stance flexion, stance extension, swing flexion, and swing extension.
+The sub-phase switching follows a finite-state machine (FSM), as shown in Fig.
+2.
 It is noted that the torque-angle relationship during the stance phase is approximately linear [25, 40].
 Inspired by this, as shown in Fig.
-2(d), this study applies a linear regression to estimate the quasi-stiffness of the torque-angle relationship, and the equilibrium angle is equal to the value of the intersection point between the regression equation and the x-axis.
-For the prosthesis control, we divide the gait phase into four sub-phases: stance flexion, stance extension, swing flexion, and swing extension.
-The sub-phase switching follows a finite-state machine (FSM), as shown in Fig.
+1(d), this study applies a linear regression to estimate the quasi-stiffness of the torque-angle relationship, and the equilibrium angle is equal to the value of the intersection point between the regression equation and the x-axis.
+While the swing phase is much more complex than the stance phase [40], we choose two linear parts that can achieve swing flexion and extension and use linear functions to approximate these two parts.
+III Experiments and Results
+III-A Experimental Setup and Protocol
+1) Prosthesis Hardware: The powered transfemoral prosthesis used in this study is shown in Fig.
 3.
-For the stance phase, the torque-angle relationship is approximately linear, which can be approximated by a linear equation.
-While the swing phase is much more complex than the stance phase [43], we choose two linear parts that can achieve swing flexion and extension and use linear functions to approximate these two parts.
-III Results
-III-A Experimental Setup
-The powered transfemoral prosthesis used in this study is shown in Fig.
-6(a).
 Both knee and ankle joints are driven by frameless motors (ILM motor kit, RoboDrive, Seefeld, Germany) with a 50:1 harmonic reducer (CSD-25-50-2A-GR, Nagano Prefecture, Japan).
 The frameless motors are actuated by drives (G-MOLTWIR50/100SEO, Elmo Motion Control, Petah Tikva, Israel).
-A 6-axis force/torque sensor (Sunrise Instruments, Nanning, China) is employed to measure the (ground reaction forces) GRFs.
-A transfemoral amputee patient is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework.
+A 6-axis force/torque sensor (Sunrise Instruments, Nanning, China) is employed to measure the GRFs (ground reaction forces).
+2) Experimental Protocol: A transfemoral amputee patient (44 years old, 60 kg, 170 cm, 24 years post-amputation) is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework.
+To investigate the task-adaptive performance of the proposed control framework, steady-state walking task trails including varying speeds (from 0.4 m/s to 1.0 m/s, m/s deviations) and varying inclines (from 2 deg to 8 deg, deg deviations) are conducted on an instrumented split-belt treadmill (Bertec, Columbus, OH, USA).
+The initial trials began with a period of acclimation in which the amputee participant first walked on the treadmill at a self-preferred speed (about 0.6 m/s) until feeling comfortable.
+After tuning and acclimation for a period of time, the participant walked on the treadmill again and performed each steady-state task trail for 30 seconds.
 The participant provided informed consent, and the Southern University of Science and Technology Institutional Review Board approved the experiments.
-III-B Experimental Results of the Proposed Framework
-Due to the space limitations, this study presents an example of varying walking speed tasks to show how the learning quasi-stiffness control framework works.
+3) Benchmark FSMIC Setup: For comparison, this article introduces a benchmark FSMIC based on the existing studies [3, 17].
+The selection of this FSMIC as a benchmark stems from its extensive usage, simplicity, and capacity to reproduce human-like kinematics and kinetics after appropriate impedance tuning.
+During the benchmark FSMIC tuning period, combined with feedback from both the participants and the prosthetist, the tuning process persisted until the authors, prosthetist, and participant expressed satisfaction with the attained natural gait.
+III-B Experimental Results
+Due to the space limitations, this study presents an example of varying walking speed tasks and varying inclines walking tasks to show how the learning quasi-stiffness control framework works.
 The learning process of the GPR-based target feature estimation and KMP-based torque-angle relationship construction is offline on the joint kinematics and kinetics obtained from the open-source dataset [37].
 1) Target Features Estimation: Fig.
-2(b) shows the experimental results of target features prediction, where the blue dots denote the demonstrated target features and the purple dots represent the predicted target features of a new task (new speed).
+1(b) shows the experimental results of target features prediction, where the blue dots denote the reference target features and the purple dots represent the predicted target features of a new task (new speed).
 The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks.
@@
 4 shows the torque-angle relationship learning and reconstruction in a new task.
-The distribution of the demonstrated torque-angle relationship can be learned by GMM.
+The distribution of the human reference torque-angle relationship can be learned by GMM.
 Based on the learned GMM, the reference torque-angle relationship can be reproduced by GMR.
 As shown in Fig.
-4(a), the blue torque-angle curves denote the human demonstrations of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR.
+4(a), the blue torque-angle curves denote the human reference trajectoires of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR.
 As shown in Fig.
-4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the demonstrated torque-angle relationships (blue curves) and target features (purple dots) in a new task.
-3) Quasi-stiffness Estimation: Fig.
-5 shows using linear functions to approximate several parts of the knee and ankle torque-angle relationships in the varying speed walking tasks.
-Table I shows the estimated stiffness and equilibrium angles from the torque-angle relationships of different speed tasks and these impedance parameters are used in the powered transfemoral prosthesis experiments.
-4) Performance of the Proposed Method in Experiments: Fig.
-6(b) shows a series of photos of a transfemoral amputee performing varying speed and inclines walking tasks.
+4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the human reference torque-angle relationships (blue curves) and target features (purple dots) in a new task.
+3) Quasi-stiffness Estimation: Fig. 5 shows using linear functions to approximate several parts of the knee and ankle torque-angle relationships in the varying speed walking tasks.
+4) Amputee Participant Experimental Results: Fig.
+6 shows a series of gait trails of a transfemoral amputee performing varying speed and inclines walking tasks.
 Fig.
@@
 The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry.
-| Task (m/s) | Joint | Quasi-stiffness (Nm/deg) | Equilibrium Angle (deg) | ||||||
-| 0.4 | knee | 4.13 | 0.59 | 0.08 | 0.19 | 12 | 22 | 30 | 15 |
-| ankle | 0.82 | 6.14 | 0.53 | 0.59 | -5 | -2 | 8 | -8 | |
-| 0.6 | knee | 5.68 | 0.70 | 0.17 | 0.26 | 12 | 21 | 57 | 17 |
-| ankle | 0.76 | 5.05 | 0.53 | 0.64 | -5 | 0 | 8 | -8 | |
-| 0.8 | knee | 5.79 | 0.79 | 0.27 | 0.45 | 13 | 22 | 69 | 18 |
-| ankle | 0.78 | 4.23 | 0.45 | 0.69 | -6 | -4 | 9 | -7 | |
-| 1.0 | knee | 5.85 | 0.85 | 0.36 | 0.52 | 13 | 22 | 76 | 20 |
-| ankle | 1.03 | 3.89 | 0.36 | 0.72 | -5 | -7 | 10 | -8 | |
+For comparison, Fig.
+9 shows the observed mean kinematics and kinetics trajectories of the proposed learning quasi-stiffness control framework and the benchmark FSMIC.
+Fig.
+10 shows the RMSE of the observed kinematics and kinetics generated by the proposed method and the benchmark FSMIC with respect to the able-bodied references.
+The angle and moment RMSE of the proposed method are usually smaller than the benchmark.
 IV Discussion
-IV-A Evaluation of the Proposed Control Method
+IV-A Evaluation of the Proposed Framework
 This work proposed a learning task-adaptive quasi-stiffness controller for a powered transfemoral prosthesis that generalizes across tasks.
@@
 As shown in Fig.
-7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking speed and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles.
-The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which indicates that the quai-stiffness control has the potential to improve gait symmetry.
+7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking tasks and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles.
+As speed or incline increases, the proposed method can generate biomechanical performances similar to the intact limb of the user and healthy individuals.
+For instance, as the incline increases, ankle joint torque correspondingly increases to provide individuals with limb amputations more push-off torque.
+The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories (also the lower RMSE of the observed kinematics with respect to the intact leg kinematics shown in Fig.
+8(a)), which indicates that the quasi-stiffness control has the potential to improve gait symmetry.
 Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
+This article introduces the widely used FSMIC controller as a benchmark for comparison.
+Fig.
+9 and 10 show the proposed learning quasi-stiffness control framework meets the benchmark’s performance in most walking tasks.
+Fig.
+10(b) indicates that the proposed method surpasses the benchmark’s performance in generating the biological kinetics.
+As shown in Fig.
+9 and Fig.
+10(a), the FSMIC’s performance was not drastically worse than the proposed method in the tested metrics, it required an average of 20 min of tuning per tuned task.
+The proposed method highlights its advantages in reproducing biomimetic kinematics and kinetics similar to able-bodied references, aligns with the performance of the FSMIC without requiring manual impedance tuning, and reveals the potential to enable autonomous task adaptation during continuously varying walking tasks.
 IV-B Limitations and Future Work
-A limitation of this study is that the torque-angle relationship of the swing phase tends to be more nonlinear and complex as the task becomes more difficult (higher walking speed or more challenging terrain conditions), which means that using linear regression to estimate quasi-stiffness may introduce larger errors.
-In future work, we will investigate the use of nonlinear regression to estimate the swing phase quasi-stiffness.
+A limitation of this study is that the torque-angle relationship of the swing phase tends to be more nonlinear and complex as the task becomes more difficult (composite tasks, such as walking speed and incline changing simultaneously), which means that using linear regression to estimate quasi-stiffness may introduce more significant errors.
+In future work, we will investigate using nonlinear regression to estimate the swing phase quasi-stiffness.
 Nevertheless, this work focuses on designing a powered transfemoral prosthesis control framework that generalizes across various tasks.
 Experimental results and walking performance (see the accompanying video) preliminarily verify that the framework has the potential to achieve task-adaptive control.
-Future studies will invite more participants to verify the performance of the proposed framework in outdoor experiments.
-V Conclusion
-This study proposed a learning task-adaptive quasi-stiffness control framework for a powered transfemoral prosthesis working over variable tasks.
-Learning from the human demonstration method is developed for the torque-angle relationship reconstruction in a new task, which consists of a GPR-based target features prediction model and a KMP-based trajectories reconstruction model.
-A transfemoral amputee participant demonstrated the control framework’s ability to generalize across tasks by varying speeds and inclines walking experiments.
-The experimental results validate that the proposed control framework has the potential to reproduce biomimetic joint torques and angles for a powered prosthesis during variable-task walking.
+Future studies will focus on evaluating the performance of the proposed method in stair climbing tasks.
+Besides, we will invite more participants to verify the performance of the proposed framework and investigate the user preference of the proposed method and the FSMIC in different tasks.

### v3

@@
-Tuning-free Quasi-stiffness Control Framework of a Powered Transfemoral Prosthesis for Task-adaptive Walking
+A Learning Quasi-stiffness Control Framework of a Powered Transfemoral Prosthesis for Adaptive Speed and Incline Walking
 Abstract
-Impedance-based control represents a prevalent strategy in the development of powered transfemoral prostheses.
-However, creating a task-adaptive, tuning-free controller that effectively generalizes across diverse locomotion modes and terrain conditions continues to be a significant challenge.
-This letter proposes a tuning-free and task-adaptive quasi-stiffness control framework for powered prostheses that generalizes across various walking tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part.
+Impedance-based control represents a prevalent strategy in the powered transfemoral prostheses because of its ability to reproduce natural walking.
+However, most existing studies have developed impedance-based prosthesis controllers for specific tasks, while creating a task-adaptive controller for variable-task walking continues to be a significant challenge.
+This article proposes a task-adaptive quasi-stiffness control framework for powered prostheses that generalizes across various walking tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part.
 A Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in a new task.
@@
 Powered transfemoral prostheses hold promise in restoring the lost joint function and significantly enhancing the mobility of these individuals [1, 2].
-While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8] and controller designs [9, 10, 11, 12, 13, 14, 15, 16] of powered prostheses over the past y, developing a prosthesis control framework that generalizes across various locomotion modes and terrain conditions remains a challenge [2].
+While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8] and controller designs [9, 10, 11, 12, 13, 14, 15, 16] of powered prostheses over the past years, developing a prosthesis control framework that generalizes across various locomotion modes and terrain conditions remains a challenge [2].
 Impedance-based control is the most commonly used approach in powered prostheses because of its ability to reproduce natural walking.
@@
 Secondly, the quasi-stuffiness control strategy is introduced after obtaining the torque-angle relationship to design a task-adaptive controller that generalizes across tasks.
+This study aims to develop a task-adaptive, tuning free quasi-stiffness control framework for a powered transfemoral prosthesis that can reproduce natural walking and biomimic joint kinematics and kinetics similar to those of able-bodied people.
 The main contributions are summarized as follows:
@@
 As shown in Fig.
-1, the proposed tuning-free, task-adaptive quasi-stiffness control framework consists of a task estimation part, a torque-angle relationship reconstruction part, and a quasi-stiffness control part.
+1, the proposed task-adaptive quasi-stiffness control framework consists of a task estimation part, a torque-angle relationship reconstruction part, and a quasi-stiffness control part.
 The task estimation part perceives amputee locomotion information and terrain conditions.
@@
 These methodologies facilitate the estimation of task parameters .
+In this study, we used the speed estimation method in [19, 36] and the incline estimation method in [33, 37].
+The task estimate accuracy is shown in Table I.
+Task | Accuracy | Reference Method |
+Speed | 94 % | [19, 36] |
+Incline | 95 % | [33, 37] |
 The torque-angle relationship reconstruction part includes a GPR-based target features prediction model and a KMP-based task-specific torque-angle relationship reconstruction model.
@@
 II-B Target Features Prediction
-An open-source dataset [37] is implemented as the human reference trajectories, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights).
+Two open-source dataset [38, 39] are implemented as the human reference trajectories, which contains tasks including multiple locomotion modes (flat walking and ramp climbing) and terrain conditions (walking speeds (28 speeds ranging from 0.5 to 1.85 m/s in 0.05 m/s increments) and ramp inclinations (6 inclination angles in [38] and 4 inclination angle in [39]), rang from -10 deg to 10 deg with 5 deg increments).
+Notably, for varying walking speeds on treadmill, we divided 70 % of the data as training set and 30 % as testing set; for varying inclines, we used all the incline walking data in [38] and [39] as training set and the data we collected under the motion capture system (rang from -8 deg to 8 deg in 2 deg increments) as testing set.
 To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the reference trajectory method for target features prediction, as shown in Fig.
-1(b).
+2(b).
 Assume there are reference trajectories and define the train data as
@@
 The sub-phase switching follows a finite-state machine (FSM), as shown in Fig.
-2.
+3.
 It is noted that the torque-angle relationship during the stance phase is approximately linear [25, 42].
@@
 2) Experimental Protocol: A transfemoral amputee patient (44 years old, 60 kg, 170 cm, 24 years post-amputation) is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework.
+Notably, to evaluate the performance of the GPR model used in the target features estimation and the KMP model used in the torque-angle curve reconstruction, we used the walking speeds (from 0.2m/s to 0.8 m/s in 0.2 m/s increments) and ramp inclinations (from -8 deg to 8 deg in 2 deg increments) as test set.
 To investigate the task-adaptive performance of the proposed control framework, steady-state walking task trails including varying speeds (from 0.4 m/s to 1.0 m/s, m/s deviations) and varying inclines (from 2 deg to 8 deg, deg deviations) are conducted on an instrumented split-belt treadmill (Bertec, Columbus, OH, USA).
@@
 1) Target Features Estimation: Fig.
-1(b) shows the experimental results of target features prediction, where the blue dots denote the reference target features and the purple dots represent the predicted target features of a new task (new speed).
+2(b) shows the experimental results of target features prediction, where the blue dots denote the reference target features and the purple dots represent the predicted target features of a new task (new speed).
 The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks.
+The prediction errors were calculated by RMSE (root mean squared error) between reference and predicted target joint angles and joint torques, as shown in Fig.
+4.
 2) Torque-angle Relationship Reconstruction: Fig.
@@
 As shown in Fig.
-4(a), the blue torque-angle curves denote the human reference trajectoires of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR.
+5(a), the blue torque-angle curves denote the human reference trajectoires of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR.
 As shown in Fig.
-4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the human reference torque-angle relationships (blue curves) and target features (purple dots) in a new task.
+5(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the human reference torque-angle relationships (blue curves) and target features (purple dots) in a new task.
+The reconstruction errors were also calculated by RMSE between reference and predicted torque-angle relationship, as shown in Fig.
+6.
 3) Quasi-stiffness Estimation: Fig. 7 shows using linear functions to approximate several parts of the knee and ankle torque-angle relationships in the varying speed walking tasks.
@@
 Fig.
-7 presents that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references.
+9 and 10 present that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references.
 The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which also indicate that the quai-stiffness can improve gait symmetry.
 Fig.
-8 shows the root mean squared error (RMSE) in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments.
+11 shows the RMSE in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments.
 The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry.
@@
 As shown in Fig.
-7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking tasks and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles.
+9 and 10, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking tasks and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles.
 As speed or incline increases, the proposed method can generate biomechanical performances similar to the intact limb of the user and healthy individuals.
@@
 The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories (also the lower RMSE of the observed kinematics with respect to the intact leg kinematics shown in Fig.
-8(a)), which indicates that the quasi-stiffness control has the potential to improve gait symmetry.
+11(a)), which indicates that the quasi-stiffness control has the potential to improve gait symmetry.
 Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
@@
 Fig.
-9 and 10 show the proposed learning quasi-stiffness control framework meets the benchmark’s performance in most walking tasks.
+12 and 13 show the proposed learning quasi-stiffness control framework meets the benchmark’s performance in most walking tasks.
 Fig.
-10(b) indicates that the proposed method surpasses the benchmark’s performance in generating the biological kinetics.
+13(b) indicates that the proposed method surpasses the benchmark’s performance in generating the biological kinetics.
 As shown in Fig.
 12 and Fig.
-10(a), the FSMIC’s performance was not drastically worse than the proposed method in the tested metrics, it required an average of 20 min of tuning per tuned task.
+13(a), the FSMIC’s performance was not drastically worse than the proposed method in the tested metrics, it required an average of 20 min of tuning per tuned task.
 The proposed method highlights its advantages in reproducing biomimetic kinematics and kinetics similar to able-bodied references, aligns with the performance of the FSMIC without requiring manual impedance tuning, and reveals the potential to enable autonomous task adaptation during continuously varying walking tasks.